OpenStudy (anonymous):

Find the slope of the line containing the given pair of points. If the line is undefined, state so. Please be sure to show all of your work. (1, -9) and (-2, -3)

OpenStudy (anonymous):

is it M=(1-9)/(-2-3)=10/5

OpenStudy (anonymous):

y=(10/5)x+b -7=-4+b -7=b l y (10/5)x -7